OpenStudy (anonymous):

Determine the resultant and equilibrant of each pair of forces acting on an object: f1 has a magnitude of 5 N acting due east, and f2 has a magnitude of 12 N acting due north.

OpenStudy (anonymous):

Okay, if we plot this on an XY Graph We'll call "north" the Y-axis, and east the X-Axis. Draw an Arrow 12 units up the y axis. Draw another arrow 5 units right on the x-axis from the tip of the 12 N line. The resultant is the arrow that connects the start of the Y arrow with the end of the x arrow. A^2 + B^2 = C^3 (Pythagorean Theorem) So Resultant = \[\sqrt{5^2 + 12^2}\] Angle = \[Tan (\theta) = y / x = 5 / 12 \] Do the arc tangent to get the angle

OpenStudy (anonymous):

 You can also draw the arrows in the oppsoite order